# Quantum Complexity of Nonlocal Field Theories

### v2

@@ 1 Introduction and Summary
In Sec. 5 we discuss our results in the context of past work and provide an outlook for future work. In the Appendix A, we revisit the holographic entanglement entropy of the dipole deformed SYM theory to facilitate comparison with the subregion volume complexity, especially to determine the critical subregion length as a function of the deformation parameter and the fact that the critical subregion size turns out to the same for both observables - be it entanglement entropy or subregion volume complexity.
@@ 2 Dipole Deformed SYM theory
where sum of all the dipoles vanishes [^2] This leads to the fact that all planer contribution in the deformed theory is same as the ordinary theory. $\sum_{i}L_{i}=0=\sum_{i}p_{i}$, due to cyclicity of trace. In general, when multiple fields are involved in the product there can be ordering ambiguity, but as we are interested in gauge theories there is a natural ordering imposed by the gauge group. The requirement of associativity over the star products, constrain the dipole length associated with the star product of two fields to be sum of the dipole length of individual fields [^3] A general choice of the dipole vector which ensure this is $\vec{L}_{i}=Q_{i}\vec{L}$, where $Q_{i}$ are the conserved charges and $\vec{L}$ is a constant vector.,
@@
where the bulk background parameter $a=\lambda\tilde{L}$ is related to the length scale of nonlocality (noncommutativity) of the boundary dual field theory $L$ as $L=2\pi\tilde{L}$, with $\lambda$ being the t’Hooft coupling $\lambda=4\pi g_{YM}^{2}N$. This dual geometry is a deformation (warping) of $AdS_{5}\times S^{5}$ along one AdS direction $x$ and one $S^{5}$ direction by a factor of $f(u)$. The deformation on the $S^{5}$ is such that the radius of $S^{1}$ fibration over $C\mathbb{P}^{2}$ is modified to $f(u)$. It is evident from the metric, dilaton and B-field expressions that in the commutative regime ($a\rightarrow 0$) we recover the usual $AdS_{5}\times S^{5}$ background with vanishing dilaton and B-field. The generic features of this geometry has been studied in detail in Bergman:2001rw , and they have shown the boundary metric is degenerate and this degeneracy sources the nonlocality along the $x$-direction in the boundary theory. Also this theory is free from any UV/IR mixing, which are present in noncommutative Super-Yang-Mills theory(NCSYM), this makes it a simpler example of non-commutative theory. While the study of correlation function in this background remains an open problem, a study of Wilson loops of the dipole theories has been carried out using the gravity dual in Alishahiha:2002ex ; Huang:2007ux ; Huang:2007dg , and it has shown presence of an dipole-dipole interaction term in the interquark potential, in addition to the usual gauge theory coulomb potential. This dipole-dipole interaction produces an attractive force between gluons within a glueball and also lead to an additional confining force between quarks in baryons, which shows an indication of new baryon phase due to this dipole-dipole attraction Huang:2008uh . Various aspects of the dipole deformed gauge theory has been studied Chiou:2003sh ; Fischler:2013gsa ; Karczmarek:2013xxa over the years using holographic tools and several interesting results has been found. In the study of entanglement entropy in dipole deformed theory Karczmarek:2013xxa it has been found that entanglement entropy follows a volume law in the region smaller than the non-commutative length scale rather than an area law. This leaves the study of behavior of complexity in the dipole deformed theory as an attractive outstanding problem, which might shed more light on the phase structure of the dipole deformed gauge theory. In the remainder of this section, we investigate the complexity characteristics of the dipole deformed gauge theory via three holographic probes, namely full volume complexity, the subregion volume complexity as a function of the subregion size (in particular the dependence of the subregion complexity on the length or extension along the dipole deformation) and finally the action complexity (using the full 10d SUGRA action). Overall, we observe that while in the near commutative/ near local regime (small $a$) the complexity does reproduce local volume law scaling(s), in the strongly nonlocal regime (large $a$), complexity departs from volume law scaling with the scaling exponent being independent of the dipole deformation parameter aa! We comment on our findings at the end of the section.
@@ 2.1 Volume complexity (CV)
where, $c$ is the central charge $c=\frac{N^{2}}{4}$ in large $N$ limit and we have introduced the ”lattice spacing” regulator in the boundary $\epsilon$ via the bulk IR cutoff $u_{b}=\frac{1}{\epsilon}$.
@@ 2.2 Subregion volume complexity
Here as the functional $\mathcal{L}(u,u^{\prime})$ does not explicitly depends on the co-ordinate $x$, we have a conserved Hamiltonian,
@@
where, $u_{*}$ is the extremum values of $u$ on the RT surface, and $u_{b}$ is the boundary UV cutoff. From (16) we can replace $u^{\prime}(x)$ in terms of $u_{*}$, this simplifies the area of co-dimension two surface,
@@
A full numerical analysis of the area/entanglement entropy and a comparison of final results with Karczmarek:2013xxa has been presented in appendix A. Here we focus on the subregion complexity, and consider the volume of co-dimension one surface bounded by RT surface.
@@
We could not proceed analytically with the above integral, therefore we move forward with numerics, and analyze the behavior of subregion complexity with the finite subregion length $l$ , shown in figure 1.
@@ 2.3 Action Complexity
$S_{bdy}=-(\pm)\frac{1}{\tilde{\kappa}_{10}}\int d^{8}X\,d\lambda\,\sqrt{-\tilde{\gamma}}\,\tilde{\kappa}$ (42)
where, $(\pm)$ depends on the if volume of interest lies in future or past of the null boundary, $\gamma$ is the induced metric on the null hypersurface, $\lambda$ parameterize the null generators, and $\kappa$ is given by,
$n^{\mu}\nabla_{\mu}n^{\nu}=\kappa n^{\nu},\quad n^{\mu}=\frac{\partial x^{\mu}}{\partial\lambda}$ (43)
where, $n^{\mu}$ is null normal vector on the hypersurface, and $\kappa$ measures the failure of $\lambda$ to be an affine parameter. This boundary contribution is valid in Einstein frame, while the bulk dual of dipole theory is known in string frame. Therefore we need to transform the quantities involved in string frame through a conformal transformation. As in 10D the bulk metric in string frame and Einstein frame are related as,
@@
$S_{bdy}$ $=$ $S_{\mathcal{N^{+}}}+S_{\mathcal{N^{+}}}$ (46)
$S_{\mathcal{N^{+}}}$ $=$ $\frac{1}{\tilde{\kappa}_{10}}\int d^{8}X\,d\lambda_{+}\,\sqrt{-\tilde{\gamma}}\,\tilde{\kappa}_{+}$
$S_{\mathcal{N^{-}}}$ $=$ $-\frac{1}{\tilde{\kappa}_{10}}\int d^{8}X\,d\lambda_{-}\,\sqrt{-\tilde{\gamma}}\,\tilde{\kappa}_{-}$
We have chosen the affine parameter $\lambda_{\pm}=\mp u$, so the null normal vector is such that $n^{\mu}_{\pm}=\mp\frac{1}{R}\frac{\partial x^{\mu}}{\partial u}=\alpha_{\pm}(\frac{1}{u^{2}},\mp 1,\vec{0})$ it gives a vanishing $\kappa=0$. Therefore only contribution to the boundary action comes due to the frame change,
@@
$S_{bdy}=S_{\mathcal{N}_{+}}+S_{\mathcal{N}_{-}}=2S_{\mathcal{N}_{+}}$ (49)
taking $\alpha_{+}=\alpha_{-}=1$
$\mathcal{C}_{\mathcal{A}}^{bdy}=\frac{2W^{3}\tan^{-1}\left(au_{b}\right)}{a^{3}}-\frac{2W^{3}u_{b}}{a^{2}}+\frac{2}{3}W^{3}u_{b}^{3}$ (50)
@@
$\mathcal{C}^{joint}_{\mathcal{A}}$ $=$ $-2W^{3}u_{b}^{3}\ln\left(\frac{\alpha^{+}\alpha^{-}}{u_{b}^{2}}\right)\$(53)
This is exactly same as the joint term contribution in case of pure AdS Reynolds:2016rvl . Taking $\alpha^{+}=\alpha^{-}=1$
$\mathcal{C}_{\mathcal{A}}^{joint}=4W^{3}u_{b}^{3}\ln\left(u_{b}\right)$ (54)
LMPS counter term: To retain the action on WDW patch to remain invariant under reparameterization of parameter $\lambda$, we need to add an counter term known as LMPS counter termLehner:2016vdi given by,
$\Delta S=-(\pm)2\int d^{8}Xd\lambda\sqrt{\tilde{\gamma}}\penalty 10000\ \Theta\ln{(\Theta)},\quad\Theta=\frac{1}{\sqrt{\gamma}}\frac{\partial\sqrt{\gamma}}{\partial\lambda}$ (55)
where, the plus and minus sign is determined based on volume of interest lies in the future or past of the considered null boundary.
$\Delta S_{\mathcal{N}^{+}}$ $=$ $-2\int d^{8}Xd\lambda^{+}\sqrt{\tilde{\gamma}}\penalty 10000\ \Theta_{+}\ln{|\Theta_{+}|},\quad\Theta_{+}=\frac{1}{\sqrt{\tilde{\gamma}}}\frac{\partial\sqrt{\tilde{\gamma}}}{\partial\lambda^{+}}$
$\Delta S_{\mathcal{N}^{-}}$ $=$ $2\int d^{8}Xd\lambda^{-}\sqrt{\tilde{\gamma}}\penalty 10000\ \Theta_{-}\ln{|\Theta_{-}|},\quad\Theta_{-}=\frac{1}{\sqrt{\tilde{\gamma}}}\frac{\partial\sqrt{\tilde{\gamma}}}{\partial\lambda^{-}}$
As we have choosen the affine parameter, $\lambda^{+}=\mp u$
@@
$\Delta S$ $=$ $\Delta S_{\mathcal{N}^{+}}+\Delta S_{\mathcal{N}^{-}}$
$=$ $2\Delta S_{\mathcal{N}^{+}}$
$\mathcal{C}_{\mathcal{A}}^{LMPS}$ $=$ $\frac{4}{3}W^{3}u_{b}^{3}+4W^{3}\ln(3)u_{b}^{3}-4W^{3}u_{b}^{3}\ln\left(u_{b}\right)$ (57)
Topological counter term: According to Kurlyand:2022vzv one must add an additional term to the bulk action of type IIB SUGRA such that in case of $AdS_{5}\times S^{5}$ the reduced bulk action over $S^{5}$ gives the volume of $AdS_{5}$,
@@
$C_{\mathcal{A}}=$ $4W^{3}u_{b}^{3}\ln 3+\frac{12W^{3}\tan^{-1}\left(au_{b}\right)}{a^{3}}\left(1-\lambda^{2}\right)-\frac{12W^{3}u_{b}}{a^{2}}\left(1-\lambda^{2}\right)+4W^{3}u_{b}^{3}\left(1-\lambda^{2}\right)$ (61)
where, $\lambda$ is t’Hooft coupling. The leading term, displays a volume law scaling and is independent of the dipole-deformation parameter $a$. Next, notice that the full action complexity, coincides with the pure AdS results which matches the volume complexity upto a numerical factor, when the the t’Hooft coupling is one. This is a remarkable feature of DDSYM theory that the complexity has an explicit dependence on the ’t-Hooft coupling parameter. Perhaps this points to the validity of holographic theory even for finite t’Hooft coupling, i.e. beyond the naive supergravity regime, in contrast to the local SYM-AdS duality case. Asymptotically (large $u_{b}$),
$C_{\mathcal{A}}\sim 4u_{b}^{3}R^{3}\ln(3)-4u_{b}^{3}W^{3}\left(\lambda^{2}-1\right)+\frac{12W^{3}u_{b}}{a^{2}}\left(\lambda^{2}-1\right)-\frac{6\pi W^{3}}{a^{3}}\left(\lambda^{2}-1\right)$ (62)
@@
    $\mathcal{C}_{\mathcal{A}}=4W^{3}\ln(3)u_{b}^{3}+\frac{12}{5}a^{2}W^{3}u_{b}^{5}+O\left(a^{3}\right)$ (63)
@@
    $\mathcal{C}_{\mathcal{A}}=4W^{3}u_{b}^{3}\left(1+\ln{3}-\lambda^{2}\right)+\frac{12W^{3}\lambda^{2}u_{b}}{a^{2}}$ (64)
    The strong nonlocality regime result (64) is worth reflecting on for a moment. For complexity to be positive definite, the t’Hooft coupling $\lambda$ should satisfy the (upper) bound $\lambda^{2}<1+\ln{3}$. We speculate that this upper bound signals the breakdown of the UV-completeness of the DDSYM theory analogous to the LST case Katoch:2022hdf .

### v3

@@ 1 Introduction and Summary
In Sec. 5 we discuss our results in the context of past work, the scope and limitations of our approach, and provide an outlook for future work. In the Appendix A, we revisit the holographic entanglement entropy of the dipole deformed SYM theory to facilitate comparison with the subregion volume complexity, especially to determine the critical subregion length as a function of the deformation parameter and the fact that the critical subregion size turns out to the same for both observables - be it entanglement entropy or subregion volume complexity. In Appendix B we furnish some details regarding the numerical limitations we encountered while performing the computations for the isotropic NCSYM case using Mathematica.
@@ 2 Dipole Deformed SYM theory
where sum of all the dipoles vanishes [^2] This leads to the fact that all planar contribution in the deformed theory is same as the ordinary theory. $\sum_{i}L_{i}=0=\sum_{i}p_{i}$, due to cyclicity of trace. In general, when multiple fields are involved in the product there can be ordering ambiguity, but as we are interested in gauge theories there is a natural ordering imposed by the gauge group. The requirement of associativity over the star products, constrain the dipole length associated with the star product of two fields to be sum of the dipole length of individual fields [^3] A general choice of the dipole vector which ensure this is $\vec{L}_{i}=Q_{i}\vec{L}$, where $Q_{i}$ are the conserved charges and $\vec{L}$ is a constant vector.,
@@
where the bulk background parameter $a=s\tilde{L}$ is related to the length scale of nonlocality (noncommutativity) of the boundary dual field theory $L$ as $L=2\pi\tilde{L}$, with $s$ being the t’Hooft coupling $s=4\pi g_{YM}^{2}N$. This dual geometry is a deformation (warping) of $AdS_{5}\times S^{5}$ along one AdS direction $x$ and one $S^{5}$ direction by a factor of $f(u)$. The deformation on the $S^{5}$ is such that the radius of $S^{1}$ fibration over $C\mathbb{P}^{2}$ is modified to $f(u)$. It is evident from the metric, dilaton and B-field expressions that in the commutative regime ($a\rightarrow 0$) we recover the usual $AdS_{5}\times S^{5}$ background with vanishing dilaton and B-field. The generic features of this geometry has been studied in detail in Bergman:2001rw , and they have shown the boundary metric is degenerate and this degeneracy sources the nonlocality along the $x$-direction in the boundary theory. Also this theory is free from any UV/IR mixing, which are present in noncommutative Super-Yang-Mills theory(NCSYM), this makes it a simpler example of non-commutative theory. While the study of correlation function in this background remains an open problem, a study of Wilson loops of the dipole theories has been carried out using the gravity dual in Alishahiha:2002ex ; Huang:2007ux ; Huang:2007dg , and it has shown presence of an dipole-dipole interaction term in the interquark potential, in addition to the usual gauge theory coulomb potential. This dipole-dipole interaction produces an attractive force between gluons within a glueball and also lead to an additional confining force between quarks in baryons, which shows an indication of new baryon phase due to this dipole-dipole attraction Huang:2008uh . Various aspects of the dipole deformed gauge theory has been studied Chiou:2003sh ; Fischler:2013gsa ; Karczmarek:2013xxa over the years using holographic tools and several interesting results has been found. In the study of entanglement entropy in dipole deformed theory Karczmarek:2013xxa it has been found that entanglement entropy follows a volume law in the region smaller than the non-commutative length scale rather than an area law. This leaves the study of behavior of complexity in the dipole deformed theory as an attractive outstanding problem, which might shed more light on the phase structure of the dipole deformed gauge theory. In the remainder of this section, we investigate the complexity characteristics of the dipole deformed gauge theory via three holographic probes, namely full volume complexity, the subregion volume complexity as a function of the subregion size (in particular the dependence of the subregion complexity on the length or extension along the dipole deformation) and finally the action complexity (using the full 10d SUGRA action). Overall, we observe that while in the near commutative/ near local regime (small $a$) the complexity does reproduce local volume law scaling(s), in the strongly nonlocal regime (large $a$), complexity departs from volume law scaling with the scaling exponent being independent of the dipole deformation parameter aa! We comment on our findings at the end of the section.
@@ 2.1 Volume complexity (CV)
where, $c$ is the central charge $c=\frac{N^{2}}{4}$ in large $N$ limit and we have introduced the “lattice spacing”, $\epsilon$, as an UV regulator in the boundary via the bulk IR cutoff $u_{b}=\frac{1}{\epsilon}$.
@@ 2.2 Subregion volume complexity
From this point we can proceed in two ways: we can either directly perform the definite integral involved in the area functional it will give us area of the RT surface as a function of subregion length $\mathcal{A}(l)$, or a further simplification can be made by observing that here the functional $\mathcal{L}(u,u^{\prime})$ does not explicitly depends on the co-ordinate $x$ that means there is a conserved quantity associated with this invariance. This conserved quantity let’ say $H(u(x),u^{\prime}(x))$ can be used to replace $u^{\prime}(x)$ in terms of the extremum value of $u(x)$ which we denote by $u_{*}$ i.e. $u=u_{*}$ when $u^{\prime}(x)=0$. This replacement after a variable change from the boundary coordinate $x$ to bulk coordinate $u$ simplifies the integral involved in the area functional and leaves us with both area functional and subregion length in terms of extremum value of u, $\mathcal{A}(u_{*})$ and $l(u_{*})$. We can still study area of the RT surface as a function of subregion length $\mathcal{A}(l(u_{*})$ by varying $u_{*}$, below we proceed in this way
The conserved Hamiltonian is given by,
@@
From (16) we can replace $u^{\prime}(x)$ in terms of $u_{*}$, this simplifies the area of co-dimension two surface,
@@
The integral involved in (17) and (18) refrain us to proceed further analytically.A full numerical analysis of the area/entanglement entropy and a comparison of final results with Karczmarek:2013xxa has been presented in appendix A. Here we focus on the subregion complexity, and consider the volume of co-dimension one surface bounded by RT surface.
@@
We could not proceed analytically with the above integral as well, therefore we move forward with numerics, and analyze the parametric behavior of subregion complexity $\mathcal{V}(u_{*})$ with the finite subregion length $l(u_{*})$, shown in figure 1.
@@ 2.3 Action Complexity
$S_{bdy}=-(\pm)\frac{1}{\tilde{\kappa}_{10}}\int d^{8}X\,ds\,\sqrt{-\tilde{\gamma}}\,\tilde{\kappa}$ (42)
where, $(\pm)$ depends on the if volume of interest lies in future or past of the null boundary, $\gamma$ is the induced metric on the null hypersurface, $s$ parameterize the null generators, and $\kappa$ is given by,
$n^{\mu}\nabla_{\mu}n^{\nu}=\kappa n^{\nu},\quad n^{\mu}=\frac{\partial x^{\mu}}{\partial s}$ (43)
where, $n^{\mu}$ is null normal vector on the hypersurface, and $\kappa$ measures the failure of $s$ to be an affine parameter. This boundary contribution is valid in Einstein frame, while the bulk dual of dipole theory is known in string frame. Therefore we need to transform the quantities involved in string frame through a conformal transformation. As in 10D the bulk metric in string frame and Einstein frame are related as,
@@
$S_{bdy}$ $=$ $S_{\mathcal{N^{+}}}+S_{\mathcal{N^{-}}}$ (46)
$S_{\mathcal{N^{+}}}$ $=$ $\frac{1}{\tilde{\kappa}_{10}}\int d^{8}X\,ds_{+}\,\sqrt{-\tilde{\gamma}}\,\tilde{\kappa}_{+}$
$S_{\mathcal{N^{-}}}$ $=$ $-\frac{1}{\tilde{\kappa}_{10}}\int d^{8}X\,ds_{-}\,\sqrt{-\tilde{\gamma}}\,\tilde{\kappa}_{-}$
We have chosen the affine parameter $s_{\pm}=\mp u$, so the null normal vector is such that $n^{\mu}_{\pm}=\mp\frac{1}{R}\frac{\partial x^{\mu}}{\partial u}=\alpha_{\pm}(\frac{1}{u^{2}},\mp 1,\vec{0})$ it gives a vanishing $\kappa=0$. Therefore only contribution to the boundary action comes due to the frame change,
@@
$S_{bdy}=S_{\mathcal{N}^{+}}+S_{\mathcal{N}^{-}}=2S_{\mathcal{N}^{+}}$ (49)
taking $\alpha_{+}=\alpha_{-}=\alpha$
$\mathcal{C}_{\mathcal{A}}^{bdy}=\alpha\left(\frac{2W^{3}\tan^{-1}\left(au_{b}\right)}{a^{3}}-\frac{2W^{3}u_{b}}{a^{2}}+\frac{2}{3}W^{3}u_{b}^{3}\right)$ (50)
@@
$\mathcal{C}^{joint}_{\mathcal{A}}$ $=$ $-2W^{3}u_{b}^{3}\ln\left(\frac{\alpha_{+}\alpha_{-}}{u_{b}^{2}}\right)\$(53)
This is exactly same as the joint term contribution in case of pure AdS Reynolds:2016rvl . Again taking $\alpha_{+}=\alpha_{-}=\alpha$
$\mathcal{C}_{\mathcal{A}}^{joint}=4W^{3}u_{b}^{3}\ln\left(u_{b}\right)-4W^{3}u_{b}^{3}\ln(\alpha)$ (54)
LMPS counter term: To retain the action on WDW patch to remain invariant under reparameterization of parameter $s$, we need to add an counter term known as LMPS counter termLehner:2016vdi given by,
$\Delta S=-(\pm)2\int d^{8}Xds\sqrt{\tilde{\gamma}}\penalty 10000\ \Theta\ln{(l_{ct}\penalty 10000\ \Theta)},\quad\Theta=\frac{1}{\sqrt{\gamma}}\frac{\partial\sqrt{\gamma}}{\partial s}$ (55)
where, the plus and minus sign is determined based on volume of interest lies in the future or past of the considered null boundary, and $l_{ct}$ is an arbitrary scale.
$\Delta S_{\mathcal{N}^{+}}$ $=$ $-2\int d^{8}Xds^{+}\sqrt{\tilde{\gamma}}\penalty 10000\ \Theta_{+}\ln{|\Theta_{+}|},\quad\Theta_{+}=\frac{1}{\sqrt{\tilde{\gamma}}}\frac{\partial\sqrt{\tilde{\gamma}}}{\partial s^{+}}$
$\Delta S_{\mathcal{N}^{-}}$ $=$ $2\int d^{8}Xds^{-}\sqrt{\tilde{\gamma}}\penalty 10000\ \Theta_{-}\ln{|\Theta_{-}|},\quad\Theta_{-}=\frac{1}{\sqrt{\tilde{\gamma}}}\frac{\partial\sqrt{\tilde{\gamma}}}{\partial s^{-}}$
As we have chosen the affine parameter, $s^{\pm}=\mp u$
@@
with this,
$\mathcal{C}_{\mathcal{A}}^{LMPS}$ $=$ $\frac{4}{3}W^{3}u_{b}^{3}+4W^{3}\ln(3)u_{b}^{3}-4W^{3}u_{b}^{3}\ln\left(u_{b}\right)+4W^{3}u_{b}^{3}\ln(l_{ct})$ (57)
Topological counter term: The AdS limit of the bosonic part of the 10d type IIB action when evaluated on the vacuum $AdS_{5}\times S^{5}$ paradoxically vanish and is inconsistent with the value of the would be 5d action obtained after it had been compactified over $S^{5}$. Similar disagreement had also been noted between the on-shell values of the reduced 3d action and the 10d action found in the case of $AdS_{3}\times S^{3}\times\mathcal{T}^{4}$ background supported by a 3-form flux. A natural way to resolve this problem has been suggested in Kurlyand:2022vzv where, it is assumed that the 10d action is missing some topological boundary term that restores the equivalence of its on-shell value with that of the 5d action. The additional boundary topological term is given by,
@@
$\mathcal{C}_{\mathcal{A}}$ $=$ $\frac{2\alpha W^{3}\tan^{-1}\left(\frac{a}{\epsilon}\right)}{a^{3}}-\frac{12s^{2}W^{3}\tan^{-1}\left(\frac{a}{\epsilon}\right)}{a^{3}}+\frac{10W^{3}\tan^{-1}\left(\frac{a}{\epsilon}\right)}{a^{3}}-\frac{2\alpha W^{3}}{a^{2}\epsilon}+\frac{12s^{2}W^{3}}{a^{2}\epsilon}$ (61)
$-\frac{10W^{3}}{a^{2}\epsilon}+4\left(\frac{W}{\epsilon}\right)^{3}\log(\ell)+\frac{2}{3}\alpha\left(\frac{W}{\epsilon}\right)^{3}-4\left(\frac{W}{\epsilon}\right)^{3}\log(\alpha)-4s^{2}\left(\frac{W}{\epsilon}\right)^{3}+\frac{10}{3}\left(\frac{W}{\epsilon}\right)^{3}$
$+4\log(3)\left(\frac{W}{\epsilon}\right)^{3}$
where in the final expression we have written the UV cutoff $u_{b}$ in terms of lattice spacing $u_{b}=\frac{1}{\epsilon}$. If we fix the normalization constant to be one $(\alpha=1)$ and the counter term scale to be AdS radius which has been taken to be one $(l_{ct}=R=1)$ the expression of complexity simplifies to,
$\mathcal{C}_{\mathcal{A}}=$ $4\left(\frac{W}{\epsilon}\right)^{3}\ln 3+\frac{12W^{3}\tan^{-1}\left(\frac{a}{\epsilon}\right)}{a^{3}}\left(1-s^{2}\right)-\frac{12W^{3}}{a^{2}\epsilon}\left(1-s^{2}\right)+4\left(\frac{W}{\epsilon}\right)^{3}\left(1-s^{2}\right)$ (62)
[…]
Asymptotically (small $\epsilon$),
$C_{\mathcal{A}}\sim 4\left(\frac{W}{\epsilon}\right)^{3}\ln(3)-4\left(\frac{W}{\epsilon}\right)^{3}\left(s^{2}-1\right)+\frac{12W^{3}}{a^{2}\epsilon}\left(s^{2}-1\right)-\frac{6\pi W^{3}}{a^{3}}\left(s^{2}-1\right)$ (63)
@@
    $\mathcal{C}_{\mathcal{A}}=4\left(\frac{W}{\epsilon}\right)^{3}(\ln 3-\ln\alpha+\ln l_{ct})+O\left(a^{2}\right)$ (64)
@@
    $\mathcal{C}_{\mathcal{A}}=\frac{2}{3}\left(\frac{W}{\epsilon}\right)^{3}\left(\alpha_{+}-6\log(\alpha)-6s^{2}+6\log(l_{ct})+5+6\log(3)\right)+O\left(\frac{1}{a}\right)$ (65)
    From the limiting behavior of complexity in the weak nonlocality regime it can be seen that we get the AdS result either by fixing $\alpha=l_{ct}$ or $\alpha=1$ with $l_{ct}=R=1$. The strong nonlocality regime result (65) is also worth reflecting on for a moment. For complexity to be positive definite, the t’Hooft coupling $s$ should satisfy the (upper) bound $\frac{l_{ct}+5}{6}+\ln(3)>s^{2}$, which depends expliciltly on $l_{ct}$ and reduce to $1+\ln(3)>s^{2}$ if we choose $l_{ct}=R=1$ . We speculate that this upper bound signals the breakdown of the UV-completeness of the DDSYM theory analogous to the LST case Katoch:2022hdf .